you drive 104 miles along a scenic highway and then take a 34-mile bike ride. your driving rate is 4 times your cycling rate. suppose you have no more than a total of 4 hours for driving and cycling. Let x represent your cycling rate in miles per hour. Use a rational inequality to determine the possible values of x.
Found 2 solutions by lynnlo, stanbon:
Answer by lynnlo(4176)
Answer by stanbon(75887)

you drive 104 miles along a scenic highway and then take a 34-mile bike ride. your driving rate is 4 times your cycling rate. suppose you have no more than a total of 4 hours for driving and cycling. Let x represent your cycling rate in miles per hour. Use a rational inequality to determine the possible values of x.
----
Driving DATA:
distance = 104 miles ; rate = 4x mph ; time = d/r = 104/(4x) hrs.
------
Cycling DATA:
distance = 34 miles ; rate = x mph ; time = d/r = 34/x hrs.
-----------
Inequality:
104/(4x) + 34/x <= 4 hrs
-----
Multiply thru by 4x to get:
------
104 + 34*4 <= 16x
26 + 34 <= 4x
60 <= 4x
x >= 15 mph (cycling rate)
----------------
Cheers,
Stan H.
================